OpenStudy (firejay5):

Solve each system of equation by elimination: Show work and steps. 32. 3x - 2y = -3 3x + y = 3

OpenStudy (anonymous):

wow...you still didn't get the concept...this is the 3rd problem of the same type u r posting

OpenStudy (firejay5):

I get it

OpenStudy (anonymous):

Then y r u asking for help again?

OpenStudy (firejay5):

cause I can

hero (hero):

3x - 2y = -3 3x + y = 3 3x = 3x 2y - 3 = 3 - y 2y + y = 3 + 3 3y = 6 y = 2 3x + 2 = 3 3x = 3 - 2 3x = 1 x = 1/3 (x,y) = (1/3, 2)

OpenStudy (firejay5):

@Hero You have to get rid of x, y, or both variables

hero (hero):

I did. I eliminated y

OpenStudy (firejay5):

@Hero 3x = 3x???

hero (hero):

Well yes because 3(1/3) = 3(1/3) 1 = 1

OpenStudy (firejay5):

well it confused me, because not all work was shown

hero (hero):

3x - 2y = -3 3x + y = 3 3x = 2y - 3 3x = 3 - y 3x = 3x 2y - 3 = 3 - y 2y + y = 3 + 3 3y = 6 y = 2 3x + 2 = 3 3x = 3 - 2 3x = 1 x = 1/3 (x,y) = (1/3, 2)

hero (hero):

If you don't understand that, I can't help you

OpenStudy (firejay5):

I mean we got the same answer

OpenStudy (anonymous):

Are you following Hero's solution? It's hard to tell if you are confused or are actually fine and just not getting what Hero is saying. No need for more explanation if you "get it"... but if not, clarify what you are asking at this point.

OpenStudy (firejay5):

I get it, all the work wasn't shown out is what confused